A plan of a house is drawn to a scale of 1:50. The lounge is 8cm long and 6.8cm wide. Find in centimeters, its length and it's width on the plan.


1
Expert's answer
2021-11-22T16:21:17-0500

Length: "L=8*50=400\;cm=4\;m."

Width: "W=6.8*50=340 \;cm=3.4\;m."